OpenStudy (lgbasallote):

Note: This is NOT a question. This is a tutorial. How to solve non-exact differential equations? Suppose you are given \((4xy + 3y^2 - x)\text dx + x(x+2y)\text dy = 0\). See comment below to see how!

OpenStudy (lgbasallote):

The first step is to put the Differential Equation in the form Mdx + Ndy = 0 where M and N are functions of x and/or y. In this problem, the equation is already in this form. M is a function of x and/or y which is \(4xy + 3y^2 - x\) and N is also a function of x and/or y but we need to simplify it and turn it into \((x^2 + 2xy)dy\) for convenience purposes. The next step is to find the partial derivative of M with respect to y, and the partial derivative of N with respect to x. In this problem, we take \(\frac{\partial (4xy + 3y^2 - x)}{\partial y}\) and \(\frac{\partial (x^2 + 2x)}{\partial x}\) That will give us \(4x + 6y\) for \(\frac{\partial M}{\partial y}\) and \(2x + 2y\) for \(\frac{\partial N}{\partial x}\). As you notice, \(\frac{\partial M}{\partial y} \ne \frac{\partial N}{\partial x}\). Therefore, it is NOT an exact differential equation, thus, we introduce an integrating factor to make it exact. But first, a step must be made before using the Integrating Factor formula. What is this step? There are two cases: Case I \(\frac{1}{N} (\frac{\partial M}{\partial y} - \frac{\partial N}{\partial x}) = f(x)\) Case II \(\frac{1}{M} (\frac{\partial M}{\partial y} - \frac{\partial N}{\partial x}) = g(y)\) What do these mean? Basically, you first subtract \(\frac{\partial N}{\partial x}\) from \(\frac{\partial M}{\partial y}\). Then, you simplify and factor out the difference. Then compare it with M and N. If you divide the difference by N, the quotient MUST be a function of x and x alone. If the quotient does not result in a function of x, then you divide by M, and the quotient MUST be a function of y and y alone. Let's look at our problem \[\frac{\partial M}{\partial y} - \frac{\partial N}{\partial x} = 4x + 6y - (2x + 2y) = 4x + 6y - 2x - 2y = 2x + 4y\]let's factor out 2. This gives us \(2(x+2y)\). Now, let's compare with M and N. If we divide this by M, we would be getting a function of x AND y, so we cannot use it. Therefore, we divide by N, which is \(x^2 + 2xy\). \[\frac{2(x+2y)}{x^2 + 2xy} = \frac{2\cancel{(x+2y)}}{x\cancel{(x+2y)}} = \frac{2}{x}\] this is clearly a function of x and x alone. Therefore, now we can use the formula for Integrating Factor which is \(e^{\int f(x)dx}\) or \(e^{-\int g(y)dy}\), depending on the quotient we got from the previous step. Since our quotient in the previous step was a function of x, we use the first formula. \[e^{\int f(x)dx} = e^{\int \frac{2}{x} dx} = e^{2\int \frac{1}{x}dx} = e^{2\ln x} = e^{\ln x^2} = x^2\] Therefore, \(x^2\) is our Integrating Factor. What do we do with it? We multiply it by M and N, and \(\textit{should}\) make it an exact differential equation. \[x^2(4xy + 3y^2 - x)dx + x^3 (x+2y)dy = 0\] Let's distribute. \[(4x^3 y + 3x^2 y^2 - x^3)dx + (x^4 + 2x^3 y)dy = 0\] Now, we take the partial derivative of M with respect to y, and the partial derivative of N with respect to x. \[\frac{\partial M}{\partial y} = 4x^3 + 6x^2 y \] \[\frac{\partial N}{\partial x} = 4x^3 + 6x^2y\] Therefore, \(\frac{\partial M}{\partial y} = \frac{\partial N}{\partial x}\) thus, it is now an exact Differential Equation!! So, now, we just proceed on solving the new equation via the Exact Differential Equation solution approach.

OpenStudy (turingtest):

I had no idea. Thanks, for reminding me why I like OS so much :D

OpenStudy (anonymous):

Nice of you to do these tutorials! :) You're really good with LaTeX! (I seen the color latex that you did the other day...)

OpenStudy (experimentx):

And if the IF is function of both x and y http://math.stackexchange.com/questions/37641/finding-integrating-factor-when-if-will-be-a-function-of-x-and-y

OpenStudy (experimentx):

that's a huge trouble ... i guess

mathslover (mathslover):

@lgbasallote must appreciate you ,, great work I prefer to say u as Great work SIR :) gr8 and thanks again . these type of teachers are needed in os :) keep it up SIR

OpenStudy (lgbasallote):

give me some topics and i will keep it up :p lol

mathslover (mathslover):

I will request you to post of these topics : 1) complex numbers 2) kinematics in 1g , 2g , 3g 3) circular motion ( dynamics and kinematics ) 4) Limits and differential calculus :-P

OpenStudy (lgbasallote):

i have a derivative tutorial here http://openstudy.com/study#/updates/4f9cb59de4b000ae9ed1a3e8 :D

OpenStudy (lgbasallote):

number 2 and 3 are physics though -__-

mathslover (mathslover):

:D that is why i typed there :-P i was joking :-P

OpenStudy (anonymous):

Excellent tutorial. I have one question: where doe the e^∫f(x)dx come from?

OpenStudy (experimentx):

assume IF is only function of y or x, use the condition of exactness ... you will get linear DE ... whose solution is that factor.

OpenStudy (anonymous):

Thanks @lgbasallote , love me some ODEs.

OpenStudy (anonymous):

To add to mathslover's suggestions, complex calculus might be an interesting one (how feasible fitting it into a single tutorial I don't know, however). Great tutorials you're doing.

OpenStudy (anonymous):

Complex integration only might be feasible to fit within one post.

OpenStudy (anonymous):

I don't think it's possible at all to fit conformal mapping and analytics within the same post.

OpenStudy (anonymous):

Actually, I'm reviewing my real analysis textbook (at a leisurely pace, since it's my favorite math subject), and I'm willing to create these DIY math posts concurrent to me reading. If anyone's interested... This helps me remember and learn, too. It might start off slow, though. My real analysis text covers some basic theory first, and I like reading that stuff.

OpenStudy (anonymous):

I, for one, may be interested. I've got to have something interesting to do in between the near-perpetual revising for my history test.

OpenStudy (turingtest):

@henpen here is a derivation of the expression for the integrating factor in a linear DE, I'm not sure how that proof might extend to this case though. Might be a good exercise to unite the two.

OpenStudy (turingtest):

http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

You are so amazingly smart! I envy yew ;D

OpenStudy (lgbasallote):

oh stop it you :""> im not *that* smart :">

OpenStudy (anonymous):

Excellent in terms of fundamenetals , :D For those experienced with differential equations, the integrating factor x^2 is observable ...

OpenStudy (nottim):

goddarn you math people.